# Simulator Ensembles for Trustworthy Autonomous Driving Testing

### v1

@@ 5.4.1 Simulation Enviroments
We have selected the following simulators for our study: BeamNG, Udacity and Donkey. We have chosen such simulators for the following reasons: 1) they seamlessly integrate with our lane-keeping ADAS case study, as the simulation environment, i.e., asphalt road on green grass and sunny weather, has been aligned matteoMaxibon2024 to be able to replicate the same simulation conditions across different simulators; 2) they are complementary as they differ in terms of the physics implementation or the rendering engine; 3) they are open-source or available under permissive academic-licenses (BeamNG). A brief description of the simulators provided is as follows:
@@
In our experiments, we execute MultiSim in three different configurations: in the configuration BD we use BeamNG and Donkey as simulator ensemble, in the configuration BU we use BeamNG and Udacity, and lastly in the configuration UD we use Udacity and Donkey. We also evaluate DSS in such configurations; in the tables and figures below, we prefix such configurations with DSS- (e.g., DSS-BU indicates the configuration BU with the DSS approach, while BU indicates the same configuration with the MultiSim approach). It is noteworthy that in the original DSS paper matteoMaxibon2024 only the configuration DSS-BU was proposed and evaluated. The extension of the approach to other simulators constitutes a novel experimental contribution of our work.
5.4.2 Study Subject
@@
We evaluate our approach on the DAVE-2 lane-keeping ADAS. DAVE-2 is a convolutional neural network developed for multi-output regression tasks based on imitation learning nvidia-dave2 . The model architecture includes three convolutional layers for feature extraction, followed by five fully connected layers. DAVE-2 has been extensively used in a variety of ADAS testing studies 2023-Stocco-TSE ; deeptest ; 10.1145/3238147.3238187 ; matteoMaxibon2024 ; 2021-Jahangirova-ICST ; biagiola2023boundary . The model takes as input an image representing a road scene, and it is trained to predict vehicle’s actuators commands. Our implementation includes a DNN with learned lane-keeping capabilities, whereas the throttle of the vehicle is determined via a linear interpolation between the minimum and maximum speeds matteoMaxibon2024 .
We use a pre-trained DAVE-2 model from previous work matteoMaxibon2024 that has been trained on more than 60,000 images and steering commands automatically by a PID-controlled autopilot driving the car on roads with different road complexity in the three simulators used in this work.
@@ 5.4.3 Search Setup
Search Algorithm. For the test case generation with MultiSim we use NSGA-II-D, integrated in Algorithm 1. For SingleSim we also use NSGA-II-D but employ only one simulator for the evaluation of the fitness value of a test case. We use the same hyper-parameter configuration for all algorithms, such as mutation, crossover rate and population size. To select the archive threshold $\delta$ that controls which test inputs shall be included in the archive during search, we follow the guidelines of Riccio et al. 2020-Riccio-FSE and perform preliminary experiments with different thresholds, ranging from 4 to 6. We assess the diversity of the corresponding feature map after the search, by identifying how many cells of the map are covered. As the archive threshold affects the diversity of the search, we use the coverage of the feature map to identify an appropriate threshold, which is 4.5 in our case. Further, we configure a maximum search budget of six hours for all testing approaches, which proved sufficient to explore the search space and obtain an adequate number of failures.
In the DSS approach, we also use NSGA-II-D for test generation. Test cases are produced through two independent search executions, each employing a different simulator, i.e., $S_{1}$ and $S_{2}$. The test cases generated by $S_{1}$ are subsequently re-executed in $S_{2}$, and vice-versa. A test case is classified as critical only if it has been evaluated as critical by both simulators. The results from both simulators, including all generated and migrated test cases, are then aggregated to represent the final output.
Table 1: Algorithm configuration for the approaches under test MultiSim, SingleSim and DSS. All configurations are equal across methods, beside the number of fitness functions used for MultiSim.
@@
Mutation/Crossover. The mutation extent is randomly selected in the range of $[-8,8]$, while the segment length is randomly set in the range $[10,20]$.
@@ 5.5 ADAS Preliminary Evaluation
On the one hand, considering an ADAS model from previous work matteoMaxibon2024 mitigates the threats to validity of training our own model for testing. On the other hand, vision perception models like DAVE-2 are negatively affected by changes of image distributions 2024-Lambertenghi-ICST ; 2025-Lambertenghi-ICST . Thus, we assess the driving performance of the ADAS from previous work in our environment prior to the empirical evaluation. Specifically, we have defined 16 different types of roads with a road-wise maximum curvature ranging from 0 to 0.19, while the number of turns was in the range 1 to 4, and the road length was 100 mt. The evaluation allows us, on the one hand to investigate the flakiness of the simulation environments, and on the other hand to verify whether the testing setup including the hardware and software configurations is suitable for our study.
We executed the ADAS in every simulator on each of the roads for 10 times to account for the flakiness, and assessed the quality of the driving using the XTE fitness function. The results of this experiment with a detailed description of the roads used can be found in our replication package replication-package . The results show that the ADAS exhibits low XTE variation (below 0.7m) across all roads without failing, even on roads with a high number of turns and high curvature. These observations suggest that while the vehicle’s trajectories and driving styles are not harmonized across simulators, the ADAS’s driving performance is not impacted by hardware limitations or latencies, as the variation is minimal and the ADAS does not fail. Thus, we conclude that the ADAS under test is sufficiently robust for exhaustive testing.
@@ 5.6 Implementation
All experiments have been executed on two computing devices, one PC with a Ryzen 5 5600X CPU, 32 GB RAM, and a 4090 Ti GPU, and on a Linux Laptop with an i7 CPU, 32 GB RAM, and a 790 GX GPU. To accelerate DNN computations we used CUDA 11.8 with cuDNN 8.9. Note, that due to simulator-specific requirements a considerable effort was required to identify a hardware configuration which was able to support our case study. BeamNG exhibited initially a very low FPS number when running on a device with the GPU 4090 Ti. In addition, BeamNG does not officially support the Linux OS, but runs on Windows. Udacity, in contrast, could not run on our Windows machine, so that we had to employ a second computing device, i.e., a Linux laptop, to perform the experiments for the BU configuration. Finally, we made sure that all simulators had an average of 20 FPS.
Overall, our experiments required approximately 540 hours i.e., $\approx$ 22 days (10 repetitions $\times$ 9 methods $\times$ 6 hours) for all compared test generation approaches. Additional approximately 84 hours were required to validate 304 failing tests.
@@ 5.7.1 Effectiveness (RQ1)
Figure 5: Validity rate (valid_rate) and number of valid failures (n_valid) identified for MultiSim, DSS and SingleSim averaged over 10 runs. The average validity rate is shown in each bar plot.
@@
The results for RQ1 are shown in Figure 5. The figure depicts the average number of failures and the number of valid failures, identified from 10 runs for MultiSim, SingleSim and DSS. In addition, in each bar plot the validity rate is annotated in the orange bar.
Number of (valid) failures. SingleSim identifies for all simulator configurations more failures than multi-simulators approaches. However, the highest number of valid failures is found by a SingleSim approach (i.e., SingleSim on Udacity, U) which is comparable to the number of valid failures identified by the best MultiSim approach, i.e., BD. DSS identifies, for all simulators configurations, less failures than MultiSim, which is similar to the number of valid failures. Simulator-wise, BD identifies significantly less failures than the respective single-simulator configurations, i.e., B and D, but more failures than DSS-BD. The number of valid failures is however similar for BD, B and D. The number of failures of BU is similar to the SingleSim counterparts, i.e., B and U, however BU has significantly less valid failures than B and U. The results for DSS-BU are comparable to those of BU. For UD the same considerations hold we made for BD.
@@
Validity Rate. The results in Figure 5 show that BD achieves a validity rate of 99%, which is the highest rate across all compared approaches and configurations. BU has a validity rate of 55%, which is surpassed by B, being close to that of DSS-BU. For UD the validity rate is higher compared to those of D, U, and DSS-UD.
We further analyzed the convergence of MultiSim and SingleSim by using the Hypervolume quality indicator (HV). HV is a well-applied metric in the search-based software engineering domain that measures how optimal and spread is the resulting solution set, allowing to compare Pareto-based optimization approaches LiEvaluateSBST22 ; LiQuality2019 and assess its convergence. The results, which are available in our replication package replication-package , show that HV values for SingleSim stabilize over time, while those of MultiSim achieve values close to those for SingleSim, while slightly increasing over time. This indicates that running SingleSim longer will likely not improve their effectiveness.
@@
Table 2: Effectiveness: Statistical tests (Wilcoxon and Vargha-Delaney) for the comparison between MultiSim and SingleSim, as well as between MultiSim and DSS for the metrics validity rate and number of valid failures. The letter L represents a large effect size magnitude, M represents a Medium effect, and S represents a small effect (repeated comparisons are excluded (N/A)). Effect values are annotated with a star, if the first approach of the comparison yields lower values then the compared approach.
@@
   valid_rate $\sim$0 (L) $\sim$ 0 (L) $\sim$ 0 (L) 0.01 (L) $\sim$ 0 (L) 0.72 (-) $\sim$0 (L) $\sim$0 (L)
   n_valid 1.00 (-) 0.85 (-) 0.83 (-) 0.03 (L) 0.77 (-) 0.04 (M) 0.06 (-) 0.09 (-)
@@
   valid_rate 0.11 (-) 0.77 (-) 0.43 (-) N/A 0.77 (-) 0.02 (L*) 0.77 (-) 0.56 (-)
   n_valid 0.04 (L) 0.01 (L) 0.16 (-) N/A 0.11 (-) 0.56 (-) 0.32 (-) 0.14 (-)
@@
   valid_rate 0.03 (L) 0.92 (-) 0.32 (-) N/A N/A 0.08 (-) 0.56 (-) 0.32 (-)
   n_valid 0.85 (-) 0.77 (-) 0.68 (-) N/A N/A 0.28 (-) 0.38 (-) 0.37 (-)
@@
Statistical Tests. We use the non-parametric pairwise Wilcoxon rank sum test (significance level 0.05) and the Vargha Delaney’s $A_{12}$ effect size to compare the valid_rate and the n_valid values shown in Figure 5. We adopt the following standard classification for effect size values: an effect size $e$ is small, when 0.56 $\leq$ $e$ $<$ 0.64 or 0.36 $<$ $e$leq 0.44, it is medium when 0.64 $\leq$ $e$ $<$ 0.71 or 0.29 $<$ $e$leq 0.36 and large when $e$geq 0.71 or $e$leq 0.29. Otherwise, the effect size is negligible. The results are shown in Table 2.
@@
The statistical test results show that BD yields significantly higher validity rates with large effect sizes than SingleSim, MultiSim and DSS approaches DSS-BU and DSS-UD. As for BU, the validity rate results for MultiSim are in general not better or worse than SingleSim or MultiSim. For UD, the results are only significant compared to SingleSim with Udacity (i.e., U). Regarding the number of valid failures, we see that BD outperforms DSS-BD with a medium effect size, while having a similar validity rate. Regarding BU, SingleSim B and U are significantly better than BU with a large effect size, what is consistent with the results in Figure 5. Concerning UD, we do not observe a statistical significant difference w.r.t. SingleSim and DSS.
$RQ_{1}$ (effectiveness). MultiSim identifies, on average, over all simulator configurations, 70% valid failures, followed by DSS (65%) and SingleSim (47%). In particular, the multi-simulator configuration BD (BeamNG and Donkey) significantly outperforms the configurations UD (Udacity and Donkey) and UB (Udacity and BeamNG) achieving an average validity rate of 99% failures and identifying 23 simulator-agnostic failures. BD outperforms SingleSim and multi-simulator based testing approaches in all configurations beside DSS-BD, where it achieves a comparable validity rate but with a significantly higher number of failures.

### v2

@@ 5.4.1 Simulation Environments
We have selected the following simulators for our study: BeamNG, Udacity and Donkey Car. We have chosen such simulators for the following reasons: 1) they seamlessly integrate with our lane-keeping ADAS case study, as the simulation environment, i.e., asphalt road on green grass and sunny weather, has been aligned matteoMaxibon2024 to be able to replicate the same simulation conditions across different simulators (see Figure 1); 2) they are complementary as they differ in terms of the physics implementation or the rendering engine; 3) they are open-source or available under permissive academic-licenses (BeamNG). A brief description of the simulators provided is as follows:
  • BeamNG is based on Unreal engine, implements soft-body physics, serves as the simulator in the SBFT competition gambi22sbft; liu23sbft; matteo24sbft, and it has been used extensively in the literature to benchmark testing approaches beamng; zohdinasab2021deephyperion; 2020-Riccio-FSE; Humeniuk24Reinforcement.
  • The Udacity simulator is a cross-platform and rigid-body physics developed with Unity3D udacity-simulator. While originally developed for educational purposes, it has been widely applied for ADAS testing matteoMaxibon2024; 2024-Grewal-ICST; 2022-Stocco-ASE; 2021-Stocco-JSEP; 2021-Jahangirova-ICST; DeepGuard; 2025-Baresi-ICSE.
  • The Donkey Car simulator (Donkey, hereafter) is part of the Donkey Car framework that is used to build and test small-scale self-driving physical vehicles donkeycar. Similarly to Udacity, it implements rigid-body physics, and is developed with Unity 3D. The simulator is actively used in the literature for the validation of testing approaches matteoMaxibon2024; 2023-Stocco-TSE; 2023-Stocco-EMSE; survey-small-scale-2; survey-small-scale-1; 2025-Lambertenghi-ASE.
@@
Unlike previous work matteoMaxibon2024, we make no assumptions about simulator fidelity and treat all simulators as black-box environments. In our experiments, we execute MultiSim in three different configurations: in the configuration BD we use BeamNG and Donkey as simulator ensemble, in the configuration BU we use BeamNG and Udacity, and lastly in the configuration UD we use Udacity and Donkey. We also evaluate DSS in such configurations; in the tables and figures, we prefix such configurations with DSS- (e.g., DSS-BU indicates the configuration BU with the DSS approach, while BU indicates the same configuration with the MultiSim approach). It is noteworthy that in the original DSS paper matteoMaxibon2024 only the configuration DSS-BU was proposed and evaluated. The extension of the approach to other simulators constitutes a novel experimental contribution of our work.
5.4.2 Study Subjects
To assess the generality and robustness of our approach, we evaluate it on three lane-keeping ADAS that differ in their design paradigms, from convolutional models, to transformer-based architectures leveraging global attention, and hybrid vision–trajectory policies that integrate planning information, described next:
@@ 5.4.2 Study Subjects
  • CNN-based ADAS (DAVE-2): DAVE-2 is a convolutional neural network developed for multi-output regression tasks based on imitation learning nvidia-dave2. The model architecture includes three convolutional layers for feature extraction, followed by five fully connected layers. The model take as input an image representing a road scene, and it is trained to predict vehicle’s actuators commands. We use a pre-trained DAVE-2 model from previous work matteoMaxibon2024 that has been trained on more than 60,000 images and steering commands automatically by a PID-controlled autopilot driving the car on roads with different road complexity in the three simulators used in this work. DAVE-2 has been extensively used in a variety of ADAS testing studies 2023-Stocco-TSE; deeptest; 10.1145/3238147.3238187; matteoMaxibon2024; 2021-Jahangirova-ICST; biagiola2023boundary; 2024-Lambertenghi-ICST; 2025-Lambertenghi-ICST.
  • Vision Transformer-based ADAS (ViT): ViT vit is a DNN based on the transformer architecture to predict steering commands. While Transformers were originally developed for natural language generation and understanding, they have since been widely adopted for visual tasks RangeViT, including lane-keeping, where ViT achieves results competitive with systems such as DAVE-2 2025-Baresi-ICSE. ViT receives an image of the road scene and predicts the steering command. Similarly to DAVE-2, we trained the ViT model using images and steering commands collected by the PID controller driving the car on 100 tracks for each simulator. Differently from DAVE-2, we trained ViT using approximately twice as many images (i.e., 120,000).
  • Trajectory-Control Policy ADAS (TCP): TCP tcp is a hybrid deep neural network that combines vision-based perception with trajectory-conditioned control to predict steering and throttle commands. The model adopts a two-branch architecture: a CNN-based vision encoder that extracts high-level spatial features from the input image, and a trajectory encoder (an MLP) that processes the planned waypoints provided by the controller. The two representations are fused in a joint latent space and passed through fully connected layers to produce continuous control outputs. Training follows a procedure similar to DAVE-2 and ViT. We collected synchronized driving data from 40 tracks per environment by running a PID controller to generate reference control signals. Each sample includes the RGB scene image, the corresponding trajectory points, and the ground-truth steering and throttle commands. The final dataset comprises over 62,000 samples, enabling TCP to learn a context-aware mapping between the visual scene, planned path, and control actions.
@@ 5.4.3 Search Setup
Search Algorithm. For the test case generation with MultiSim we use Algorithm 1. For SingleSim, we employ only one simulator for the evaluation of the fitness value of a test case. We use the same hyper-parameter configuration for all algorithms, such as mutation, crossover rate and population size. To select the archive threshold $\delta$ that controls which test inputs shall be included in the archive during search, we follow the guidelines of Riccio et al. 2020-Riccio-FSE and perform preliminary experiments with different thresholds, ranging from 4 to 6. We assess the diversity of the corresponding feature map after the search, by identifying how many cells of the map are covered. As the archive threshold affects the diversity of the search, we use the coverage of the feature map to identify an appropriate threshold, which is 4.5 in our case. Further, we configure a maximum search budget of six hours for all testing approaches with DAVE-2 and ViT and 3.5 hours with TCP, which proved sufficient to explore the search space and obtain an adequate number of failures.
In the DSS approach, we also use Algorithm 1 for test generation. Test cases are produced through two independent search executions, each employing a different simulator, i.e., $S_{1}$ and $S_{2}$. The test cases generated by $S_{1}$ are subsequently re-executed in $S_{2}$, and vice-versa. A test case is classified as critical only if it has been evaluated as critical by both simulators. The results from both simulators, including all generated and migrated test cases, are then aggregated to represent the final output.
Table 1: Test generation configuration for the approaches under test MultiSim, SingleSim and DSS. All configurations are equal across methods, beside the number of fitness functions used for MultiSim.
@@
Mutation/Crossover. The mutation extent is randomly selected in the range of [8,8][-8,8], while the segment length is randomly set in the range $[10,20]$.
@@ 5.5 ADAS Preliminary Evaluation
For DAVE-2, we considered an ADAS model from previous work matteoMaxibon2024, to mitigate the threats to validity of training our own model for testing. Vision perception models like DAVE-2, are negatively affected by changes of image distributions 2024-Lambertenghi-ICST; 2025-Lambertenghi-ICST; 2025-Lambertenghi-ASE. Thus, we assess the driving performance of the ADAS from previous work in our environment prior to the empirical evaluation. Specifically, we have defined 16 different types of roads with a road-wise maximum curvature ranging from 0 to 0.19, while the number of turns was in the range 1 to 4, and the road length was 100 mt. The evaluation allows us, on the one hand to investigate the flakiness of the simulation environments, and on the other hand to verify whether the testing setup including the hardware and software configurations is suitable for our study.
For ViT, we selected an existing model architecture 2025-Baresi-ICSE and trained the model on collected images with corresponding PID-based steering commands by letting the model drive on 100 diverse automatically generated roads.
For TCP, we reused the model architecture from the original work tcp, which includes a ResNet model for visual feature extraction trained in the CARLA simulator. Because the pre-trained ADAS exhibited consistent failures across our simulators, we retrained it on a unified dataset collected from all three environments through PID-controlled driving over 40 randomly generated tracks, ensuring balanced exposure to the visual and physical characteristics of each simulator. The collected data was composed of images, steering commands, vehicle position information, and target waypoint positions. The TCP model defines a parameter alpha that specifies the weight of the trajectory branch vs. the predict control branch for combined steering value prediction. We selected the alpha score, i.e., 0.1, based on preliminary experiments using validation roads. We executed all ADAS in every simulator on each of the predefined roads for 10 times to account for the flakiness, and assessed the quality of the driving using the XTE fitness function.
The results show that all ADAS exhibit low XTE variation (below 0.7m) across all roads without failing, even on roads with a high number of turns and high curvature. These observations suggest that while the vehicle’s trajectories and driving styles are not harmonized across simulators, the ADAS’s driving performance is not impacted by hardware limitations or latencies, as the variation is minimal and the ADAS does not fail. Thus, we conclude that the evaluated ADAS under test are sufficiently robust for exhaustive testing.
The results of all preliminary experiments with a detailed description of the roads used can be found in our replication package replication-package.
@@ 5.6 Implementation
All experiments have been executed on two computing devices, one PC with a Ryzen 5 5600X CPU, 32 GB RAM, and a 4090 Ti GPU, and on a Linux Laptop with an i7 CPU, 32 GB RAM, and a 790 GX GPU. To accelerate DNN computations we used CUDA 11.8 with cuDNN 8.9. Note, that due to simulator-specific requirements a considerable effort was required to identify a hardware configuration which was able to support our case study. BeamNG exhibited initially a very low FPS number when running on a device with the GPU 4090 Ti. In addition, BeamNG does not support the Linux, but runs on Windows. Udacity, in contrast, could not run on our Windows machine, so that we had to employ a second computing device, i.e., a Linux laptop, to perform the experiments for the BU/DSS-BU configuration. Finally, we made sure that all simulators run at 20 FPS.
Overall, our experiments required approximately 1,395 hours i.e., $\approx$ 58 days (3 systems $\times$ 6/10 repetitions $\times$ 9 methods $\times$ 3.5/6 hours) for all compared test generation approaches. Additional more than 25 hours were required to validate more then 6,000 failing tests.
@@ 5.7.1 Effectiveness (RQ1)
Figure 5: Validity rate (valid_rate) and number of valid failures (n_valid) identified for MultiSim, DSS, and SingleSim averaged for DAVE-2, ViT, and TCP. The average validity rate is shown in each bar plot.
@@
The results for RQ1 are shown in Figure 5. The figure depicts the average number of failures and the number of valid failures, identified from 10 runs for MultiSim, SingleSim and DSS for the system DAVE-2, and from 6 runs for the systems ViT and TCP. In addition, in each bar plot the validity rate is annotated in the orange bar.
Across all configurations and system under tests, single-simulator (SingleSim) approaches generally detect more total failures than multi-simulator ones, but many of these are invalid. The highest number of valid failures is achieved by SingleSim on Udacity (U), which is comparable to the best multi-simulator configuration, BD for DAVE-2.
While DSS yields in general fewer total failures overall, the number of valid failures remains for DAVE-2 similar to MultiSim. However, for ViT and TCP the number of valid failures is in general lower for DSS.
Simulator-wise, for DAVE-2 and ViT, BD produces fewer total failures than its single-simulator counterparts (B and D) but maintains a comparable number of valid ones. In contrast, BU identifies a similar total number of failures as B and U but significantly fewer valid cases, a pattern also observed for DSS-BU and UD. For TCP, the configurations BU and UD reach a significantly higher number of valid failures compared the single simulator-based counterparts, as well as to DSS and the multi-simulator configuration BD.
@@
The results in Figure 5 show that for DAVE-2 and ViT, BD achieves the highest validity rate across all compared approaches and configurations, (99% for DAVE-2, 98% for ViT). For DAVE-2, BU has a validity rate of 55%, which is surpassed by B, being close to that of DSS-BU. For UD the validity rate is higher compared to those of D, U, and DSS-UD.
For TCP, BU has the highest validity rate (80%), which is higher then the corresponding DSS based counterpart. Also, we can observe that all SingleSim approaches achieve lower validity rates then the worst MultiSim combination BD (0.25).
We further evaluated the convergence behavior of MultiSim and SingleSim using the Hypervolume (HV) quality indicator, a widely adopted metric in search-based software engineering for assessing both the optimality and spread of Pareto fronts LiEvaluateSBST22; LiQuality2019. The HV analysis (available in our replication package replication-package) shows that SingleSim configurations quickly reach stable HV values, whereas MultiSim achieves comparable values that continue to improve slightly over time. This suggests that extending the execution time of SingleSim would likely not yield further gains in effectiveness, while MultiSim maintains gradual improvements as the search progresses.
@@
Table 2: Effectiveness: Statistical tests (Wilcoxon and Vargha-Delaney) for the comparison between MultiSim and SingleSim, as well as between MultiSim and DSS for the metrics validity rate and number of valid failures. The letter L represents a large effect size magnitude, M represents a medium effect, and S represents a small effect (repeated comparisons are excluded (N/A)). Effect values are annotated with a star if the first approach of the comparison yields lower values than the compared approach.
@@
       valid_rate $\sim$0 (L) $\sim$ 0 (L) $\sim$ 0 (L) 0.01 (L) $\sim$ 0 (L) 0.03 (L) 0.03 (L) 0.14 (-)
       n_valid 1.00 (-) 0.85 (-) 0.83 (-) 0.03 (L) 0.77 (-) 0.03 (L) 0.03 (L) 0.06 (S)
@@
       valid_rate 0.11 (-) 0.77 (-) 0.43 (-) 0.77 (-) 0.02 (L) N/A 0.56 (L*) 0.77 (-) (-)
       n_valid 0.04 (L) 0.01 (L) N/A 0.16 (-) 0.11 (-) 1.00 (M) 0.31 (M) 0.08 (-)
@@
       valid_rate 0.03 (L) 0.92 (-) 0.32 (-) N/A N/A 0.16 (M) 0.03 (L) 0.14 (-)
       n_valid 0.85 (-) 0.77 (-) 0.68 (-) N/A N/A 0.16 (L) 0.09 (M) 0.44 (S)
ViT
BD
       valid_rate 0.03 (L) 0.03 (L) 0.03 (L) 0.03 (L) 0.59 (S) 0.04 (L) 0.03 (L) 0.14 (-)
       n_valid 0.09 (M) 0.03 (L) 0.56 (S) 0.06 (L) 0.31 (M) 0.03 (L) 0.03 (L) 0.06 (S)
UD
       valid_rate 0.03 (L) 0.03 (L) 0.03 (L) 0.03 (L) N/A 0.16 (M) 0.03 (L) 0.14 (-)
       n_valid 0.22 (M) 0.08 (M) 0.31 (M) 0.16 (M) N/A 0.16 (L) 0.09 (M) 0.44 (S)
BU
       valid_rate 0.84 (-) 0.03 (S) 0.03 (L*) N/A N/A 1.00 (M) 0.31 (M) 0.06 (-)
       n_valid 0.44 (S) 0.31 (M) 0.04 (L*) N/A N/A 1.00 (M) 0.50 (-) 0.08 (-)
TCP
BD
       valid_rate 0.04 (L) 0.69 (M) 0.04 (L) 0.56 (S*) 0.06 (L*) 0.04 (L) 0.14 (L) 0.08 (L*)
       n_valid 0.78 (L) 0.59 (S) 1.00 (-) 0.09 (L*) 0.06 (L*) 0.07 (L*) 0.04 (L) 0.36 (M*)
UD
       valid_rate 0.06 (L*) 0.16 (L*) 0.06 (L*) 0.09 (L*) N/A 0.09 (L) 0.04 (L) 0.56 (S)
       n_valid 0.09 (L*) 0.06 (L*) 0.22 (L*) 1.00 (-) N/A 0.06 (L) 0.04 (L) 0.06 (L*)
BU
       valid_rate 0.04 (L) 0.06 (L*) 0.04 (L) N/A N/A 0.04 (L) 0.04 (L) 0.29 (L*)
       n_valid 0.04 (L) 0.06 (L*) 0.04 (L) N/A N/A 0.04 (L) 0.04 (L) 0.06 (L*)
@@
Statistical Test. To assess whether the differences observed are statistical significant, we use the non-parametric pairwise Wilcoxon rank sum test (significance level 0.05) and the Vargha Delaney’s $\hat{A}_{12}$ effect size to compare the valid_rate and the n_valid values shown in Figure 5. We adopt the following standard classification for effect size values: an effect size $e$ is small, when 0.56 $\leq$ $e$ << 0.64 or 0.36 << $e$ $\leq$ 0.44, it is medium when 0.64 $\leq$ $e$ << 0.71 or 0.29 << $e$ $\leq$ 0.36 and large when $e$ $\geq$ 0.71 or $e$ $\leq$ 0.29. Otherwise, the effect size is negligible. The results are shown in Table 2.
@@
For DAVE-2, the statistical test results show that BD yields significantly higher validity rates with large effect sizes than SingleSim, MultiSim and DSS approaches DSS-BU and DSS-UD. As for BU, the validity rate results for MultiSim are in general not better or worse than SingleSim or MultiSim. For UD, the results are only significant compared to SingleSim with Udacity (i.e., U). Regarding the number of valid failures, we see that BD outperforms DSS-BD with a medium effect size, while having a similar validity rate. Regarding BU, SingleSim B and U are significantly better than BU with a large effect size, what is consistent with the results in Figure 5. Concerning UD, we do not observe a statistical significant difference w.r.t. SingleSim and DSS.
For ViT, similarly there is a statistically significant difference with large effect sizes for the valid rate between MultiSim-based and SingleSim based approaches as well as the corresponding DSS-BD baseline. Also for UD there is a statistically relevant difference compared to SingleSim approaches which is inline with the results with Figure 5. For the number of valid failures, we see statistically significant higher values for BD compared to the single simulator approach B as well to corresponding DSS counterparts.
For TCP, both BD and BU configurations exhibit significantly higher validity rates and numbers of valid failures than the single-simulator (U, D) and corresponding DSS configurations, with large effect sizes. UD also shows statistically significant improvements over most SingleSim configurations, though with smaller effects. Overall, results consistently confirm that combining complementary simulators enhances the discovery of valid failures across architectures, with BD emerging as the most effective configuration.
$RQ_{1}$ (effectiveness). Across all ADAS, MultiSim achieves higher validity rates and numbers of valid failures than both SingleSim and DSS (70% vs. 65% and 47%). Among multi-simulator settings, BD (BeamNG–Donkey) performs best for DAVE-2 and ViT (up to 99% validity), while BU (BeamNG–Udacity) excels for TCP with up to 80%. This indicates that the most effective simulator pairing depends on the SUT: BD benefits vision- and control-based systems, whereas BU better captures perception-driven failures. Statistical tests confirm that MultiSim significantly outperforms both SingleSim and DSS, except for DSS-BD, where it achieves comparable validity but a higher number of failures.